1. Dratif Corporation's working capital is $41,000 and its current liabilities are $112,000. The corporation's current ratio is closest to?

2. Stimac Corporation has total cash of $285,000, no marketable securities, total current receivables of $356,000, total inventory of $181,000, total prepaid expenses of $68,000, total current assets of $890,000, total current liabilities of $306,000, total stockholders’ equity of $2,514,000, total assets of $3,665,000, and total liabilities of $1,151,000. The company’s acid-test (quick) ratio is closest to?

3. Financial statements for Maraby Corporation appear below:

Maraby Corporation
Balance Sheet
December 31, Year 2 and Year 1
(dollars in thousands)
Year 2 Year 1
Current assets:
Cash and marketable securities $ 220 $ 190
Accounts receivable, net 190 160
Inventory 140 150
Prepaid expenses 70 80
Total current assets 620 580
Noncurrent assets:
Plant & equipment, net 1,180 1,150
Total assets $ 1,800 $ 1,730
Current liabilities:
Accounts payable $ 100 $ 120
Accrued liabilities 100 70
Notes payable, short term 160 160
Total current liabilities 360 350
Noncurrent liabilities:
Bonds payable 450 500
Total liabilities 810 850
Stockholders’ equity:
Common stock, $5 par 160 160
Additional paid-in capital 200 200
Retained earnings 630 520
Total stockholders’ equity 990 880
Total liabilities & stockholders’ equity $ 1,800 $ 1,730
Maraby Corporation
Income Statement
For the Year Ended December 31, Year 2
(dollars in thousands)
Sales (all on account) $ 1,960
Cost of goods sold 1,370
Gross margin 590
Selling and administrative expense 230
Net operating income 360
Interest expense 50
Net income before taxes 310
Income taxes (30%) 93
Net income $ 217

Maraby Corporation's working capital (in thousands of dollars) at the end of Year 2 was closest to?
Answer #1

1.) Current assets = working capital+Current liabilities = 41000+112000 = 153000

Current ratio = Current assets/Current liabilities = 153000/112000 = 1.37:1

2) Quick ratio = Quick assets/Current liabilities = (285000+356000)/306000 = 2.09:1

3) Working capital = Current assets-Current liabilities

= 620/360

Working capital = 1.72:1